Fanning Manufacturing Company began operations on January 1. During the year, it started and completed 1,770 units of product. The financial statements are prepared in accordance with GAAP. The company incurred the following costs: 1. Raw materials purchased and used-$3,060 2. Wages of production workers-$3,480. 3. Salaries of administrative and sales personnel-$1,990 4. Depreciation on manufacturing equipment-$5,142 5. Depreciation on administrative equipment-$1,775 Fanning sold 1,140 units of product Required a. Determine the total product cost for the year. b. Determine the total cost of the ending inventory. (Do not round intermediate calculations.) c. Determine the total of cost of goods sold. (Do not round intermediate calculations.) b. Total cost of ending inventory c. Total cost of goods sold
Answer #1

(a) Total product cost for the year :

Raw materials purchased and used = $ 3060

Wages of production workers    = $ 3480

Depreciation on manufacturing equipment = $ 5142

Total product cost = $ 11682

(b) Total cost of the ending inventory :

Cost of inventory per unit = $11682 ÷ 1770 = $ 6.60

Ending inventory in units = 1770 - 1140 = 630

Cost of ending inventory = $ 6.60 × 630 = $ 4158

(c) Total cost of goods sold :

COGS = $ 6.60 × 1140 = $ 7524